OpenStudy (anonymous):

Simplify the expression 6t^2-5.4-3.5+2.5t^2+3t

OpenStudy (texaschic101):

6t^2 - 5.4 - 3.5 + 2.5t^2 + 3t basically, all we have to do is combine like terms. Do you know what like terms are ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

example : like terms 5 + 4 = 9 6t + 4t = 10t 3x^2 + 4x^2 = 7x^2 unlike terms 5 + 2t 6 - 3t^3 4t^2 + 7t

OpenStudy (texaschic101):

so put your like terms together.. 6t^2 + 2.5t^2 = ?? -5.4 - 3.5 = ?? and don't forget about 3t...even though it does not have any like terms

OpenStudy (texaschic101):

you gonna try...or do you still have questions ?

OpenStudy (texaschic101):

if you still have questions, don't hesitate to ask :)